# Frontiers in Intelligent Colonoscopy

### v1

@@ Abstract
Colonoscopy is currently one of the most sensitive screening methods for colorectal cancer. This study investigates the frontiers of intelligent colonoscopy techniques and their prospective implications for multimodal medical applications. With this goal, we begin by assessing the current data-centric and model-centric landscapes through four tasks for colonoscopic scene perception, including classification, detection, segmentation, and vision-language understanding. This assessment enables us to identify domain-specific challenges and reveals that multimodal research in colonoscopy remains open for further exploration. To embrace the coming multimodal era, we establish three foundational initiatives: a large-scale multimodal instruction tuning dataset ColonINST, a colonoscopy-designed multimodal language model ColonGPT, and a multimodal benchmark. To facilitate ongoing monitoring of this rapidly evolving field, we provide a public website for the latest updates: https://github.com/ai4colonoscopy/IntelliScope.
@@
Colonoscopy, Survey, Vision-language, Multimodal Language Model, Medical Image, Abdomen, Healthcare AI.
1 Introduction
Despite declining colorectal cancer (CRC) rates in high-income countries, it remains the third most diagnosed cancer worldwide and is increasing in developing countries [1]. Colonoscopy, as an efficient method for CRC screening, utilises a flexible camera-equipped tube to visually examine the colon’s interior. As illustrated in Fig. (1-a), this clinical procedure also facilitates intervention with specialised instruments such as snares, forceps, and cautery devices to remove precancerous growths, such as serrated and adenomatous polyps. A recent study [2] indicates that incorporating artificial intelligence (AI) into colonoscopy reduces the miss rate of colorectal neoplasia by approximately 50% compared to traditional methods. This success motivates us to investigate the frontiers in intelligent colonoscopy.
Colonoscopy, an endoscopic optical imaging technique, usually presents visual patterns (e.g., non-uniform illumination, homogeneity) that differ from those of general-purpose imaging data, e.g., ImageNet [3], due to the complex and folded anatomy of the colon. This suggests that special methods are needed to interpret the colonoscopic data. In response, we begin with an investigation of the latest intelligent techniques for colonoscopy, assessing the current landscape to sort out domain-unique challenges and underexplored areas. Our analysis reveals that multimodal research in colonoscopy remains largely untapped. To bridge this gap, we contribute three efforts to the community, as illustrated in Fig. (1-b).
Figure 1: Introductory diagram. We depict (a) the anatomy of the large intestine (colon) within the digestive tract, the polypectomy procedure during colonoscopy examination, and the components of a colonoscope. The bottom figure (b) summarises three highlights of this study.
@@ 1 Introduction
Contribution. (a) We investigate the latest research progress in four colonoscopic scene perception tasks (refer to Fig. 2) from both data-centric and model-centric perspectives. Our investigation summarises key features of 63 datasets and 137 representative deep techniques published since 2015. Additionally, we highlight emerging trends and opportunities for future study. (b) We introduce ColonINST, a pioneering instruction tuning dataset tailored for multimodal research, aimed at instructing models to execute user-driven tasks interactively. Assembled from 19 publicly available sources, the ColonINST dataset contains 303,001 colonoscopy images across 62 sub-categories, reflecting diverse scenarios encountered in colonoscopy procedures. We expand these visual samples in two aspects. First, we leverage the multimodal AI chatbot, GPT-4V [4], to generate 128,620 medical captions. Second, we restructure 450,724 human-machine conversations for multimodal adaptation. (c) Leveraging the instruction tuning data, we build a multimodal language model, ColonGPT  to assist endoscopists through interactive dialogues. To ensure reproducibility for average community users, we implement ColonGPT in a resource-friendly way, using a 0.4B-parameter visual encoder SigLIP-SO [5] and a 1.3B-parameter lightweight language model Phi1.5 [6]. Unlike previous vision-language (VL) linking methods [7, 8, 9] that employ multilayer perceptrons to handle equally all tokens from the visual encoder, we propose a multigranularity adapter to selectively sample visual tokens according to their significance. This strategy reduces the visual tokens to $\sim$34of the original number without compromising performance, securing the top spot in our newly-created multimodal benchmark across three tasks. Importantly, our ColonGPT can be trained within five hours on four A100-40GB GPUs, facilitating rapid proof-of-concept development for subsequent research.
@@
Scope. This study differs from previous works in several aspects. Earlier surveys on traditional [10] and deep learning [11, 12, 13, 14] methods conducted before 2020 are now out of date. Although a recent study [15] explores various applications of colonoscopy, such as quality analysis and abnormality detection, it lacks numerical validation. Other benchmarks [16, 17, 18] are limited to specific narrow research subfields. By contrast, we delve into four tasks for colonoscopic scene perception and evaluate their current state to sort out key challenges and under-researched areas. Importantly, our vision goes beyond by laying the foundations for the coming era, a multimodal world. To embrace this era, we further undertake three initiatives: a multimodal instruction tuning dataset, a multimodal language model, and a multimodal benchmark for the community.
@@
Organisation. The remaining sections are structured as such: §2 provides a historical background and discusses the domain-unique challenges. §3 investigates 63 colonoscopy related datasets, followed by a survey of 137 deep models in §4. In §5, we introduce three initiatives towards the multimodal era: the creation of ColonINST, the technical details of ColonGPT, and a comparative multimodal benchmark along with ablative analyses. Finally, this paper is concluded in §6.
@@ C-E Crohn’s Disease [53] FMOLB’21 467 164 #C1 $\rightarrow$ Crohn’s lesions -
2 Background
2.1 Origin and evolution
The history of colonoscopy has two key milestones. In 1968, gastrointestinal surgeons Hiromi Shinya and William Wolff found a link between colonic polyps and intestinal tumours, but they lacked equipment to examine them. In 1969, they discovered Corning Incorporated’s optical fibres and collaborated with Olympus to create the fiberoptic colonoscope, a groundbreaking device to examine the colon and remove polyps using wire loops. The second milestone came in 1983 with the introduction of the electronic colonoscope [79], which allows visualisation of the colon on a screen and polyp removal using a polypectomy snare, enhancing detection rates and reducing bleeding. The 21st century brings the AI era, where computer-aided diagnosis systems provide greater precision and reliability in procedures [80]. This study explores the transformative impact of intelligent techniques for colonoscopy, which is a type of endoscopy [81], while other related techniques such as laparoscopy [82] are summarised in our appendix.
2.2 Intrinsic traits and domain-unique challenges
@@
3 Revisiting Colonoscopy Data
3.1 Medical data for colonoscopy
@@
$\bullet$ Detection datasets provide both categorical and localisation labels for targets of interest, such as colonic diseases [54, 57, 62], accessory instruments [42, 34, 42], polyps [48, 59, 23, 64, 72], endoscopic artefacts [45, 36], and gastrointestinal diseases [56, 43]. In addition, the organisation of competitions has accelerated growth within the colonoscopy community by establishing shared platforms for data collection and model evaluation, significantly advancing research areas such as the detection of intestinal disease [40] and polyp [61, 69, 38].
@@ 3.2 Discussion
$\bullet$ Data granularity requires improvement to better understand patient conditions and treatment efficacy. (a) More than a quarter of polyp-containing datasets provide fine-grained categorisation, often with inadequate detail. For example, BKAI-Small/-Large [46] provide two instance annotations of neoplastic and non-neoplastic polyps. ColonoscopicDS [26] categorises at the video level into hyperplasic, serrated, and adenoma lesions. SUN-database [54] provides fine-grained labels, documenting measurements of polyp size (height, width) and morphology (pedunculated, sessile, flat), along with their anatomical locations (e.g., rectum, sigmoid colon). Several data-centric areas remain under-explored, such as temporal lesion evolution recording, granularity improvement, graded severity tagging, and instance-level target annotation. (b) Furthermore, label orthogonality, an often overlooked issue, treats labels as isolated entities. Current works seldom discuss potential inter-class correlations, such as the co-occurrence of inflammatory bowel disease with erosion symptoms, Crohn’s disease with fistula complications, or colorectal cancer accompanied by bleeding. Future studies should consider causality [88] and comorbidity [89] to address these correlations effectively.
@@
TABLE II: Summary of classification models in colonoscopy. Dataset: CU=CU-ColonDB [25], CDS=ColonoscopicDS [26], Private=private data, HK= HyperKvasir [43], KC=Kvasir-Capsule [56]. Backbone: CaffeNet [97], D-121=DenseNet121 [98], R-12/-18/-50/-101=ResNet12/18/50/101 [99], ViT-S16 or ViT-B16 [100], MobV2=MobileNetV2 [101], R50-Att=ResNet50 with attention module [102], C3D [103], Inc-v3=Inceptionv3 [104], I3D [105]. “Customised” means a base network modified for the current task or a model independent of the base network choice. Head: classifier implemented by the fully connected (FC) and support vector machine (SVM) layers, or using the $\ell^{2}$ norm to measure the disparity between the input and output. Arch: the architectures shown in Fig. 3. Sup: learning strategies such as fully supervised (FS), semi-supervised (SS), unsupervised (US), and weakly supervised (WS). For simplicity, the following tables use consistent abbreviations unless specified otherwise.
@@
4 Revisiting Colonoscopy Models
In this section, we investigate 137 deep learning models for colonoscopic scene perception, sourced from leading conferences or journals published since 2015. First, three widely recognised topics are described, including 18 classification models (§4.1), 24 detection models (§4.2), and 86 segmentation models (§4.3). Their architectural designs are classified into five subtypes as presented in Fig. 3. Lastly, we discuss nine VL-related models in §4.4.
4.1 Classification models
@@
V2I-DETR [148] arXiv’24 video-to-image knowledge distillation SUN SUN R-50 OS BF#2 FS -
@@
$\bullet$ Remarks. We observe three aspects of the above classification models. (a) Novel visual backbones, like the state space model [149], remain underexplored. Furthermore, reformulating the classification paradigm within VL models, e.g., CLIP [150], can yield unexpected results. (b) Benchmarks for multicategory classification remain underexplored. The Kvasir series [28, 43, 56] offers valuable sources for further study. We will explore these public data on their potential synergy in §5.1. (c) Several new task settings have emerged in colonoscopy. For example, Tian et al.[59] recognise abnormal frames from colonoscopy videos from an out-of-distribution view. DAFON [115] solve an open-set classification problem within a few-shot framework.
4.2 Detection models
@@
U-KAN [236] arXiv’24 U-shaped Kolmogorov-Arnold network [237] C6 C6 Customised MF#1 - FS Link
@@
TABLE V: Comparison of image polyp segmentation models. They are evaluated using the mean scores (%) of structure measure ($\mathcal{S}$ [258]) and Dice coefficient ($\mathcal{D}$) on two test sets, with boxplots illustrating the distribution of their consistency and variability across test cases.
@@
4.3 Segmentation models
Compared to the above two topics, the segmentation research appears to be well established. This is evidenced by the extensive amount of research documented in Tab. IV.
@@ 4.3 Segmentation models
$\bullet$ Remarks. To reflect current field progress, we analyse 14 open-source image segmentation models on two popular test datasets, as shown in Tab. V. All models are trained on the same dataset from Fan et al.’s benchmark [175]. First, we observe that current learning strategies are underexplored, as evidenced by BoxPolyp, a weakly supervised model obtaining the best $\mathcal{D}$ score (91.84%) on Kvasir-SEG. In addition, some models achieve better performance, yet exhibit wider interquartile ranges in boxplots, indicating their variability in predictions. For example, Polyp-PVT, which ranks highest in the $\mathcal{S}$ score on Kvasir-SEG, shows a wider interquartile range than other models like SSFormer. From these results, we suggest several promising opportunities for future study. (a) The current gold benchmark [175] comprises less than 1.5K samples and is focused on one category (polyp). In general, scaling up both the data size and diversity could be a natural way to improve robustness and generalisability. This demand is driving innovations, such as developing a semi-auto image annotator [166, 165] to reduce expert labour and synthesising high-fidelity content via diffusion [261] and autoregressive [262] techniques. (b) Moreover, infinite data scaling is not sustainable. Developing data-efficient strategies [94, 95, 96] that require fewer or weaker labels is more cost-effective for average users in the community. (c) Finally, providing procedural support to physicians is essential, including anomaly detection, navigation planning, risk assessment, and intervention advice. We can adopt innovations from similar topics [263].
[…]
Figure 4: Details of the established ColonINST. (a) Three sequential steps to create the instruction tuning dataset for multimodal research. (b) Numbers of colonoscopy images designated for training, validation, and testing purposes. (c) Data taxonomy of three-level categories. (d) A word cloud of the category distribution by name size. (e) Caption generation pipeline using the VL prompting mode of GPT-4V [4]. (f) Numbers of human-machine dialogues created for four downstream tasks.
4.4 VL models

### v2

@@ Abstract
Colonoscopy is currently one of the most sensitive screening methods for colorectal cancer. This study investigates the frontiers of intelligent colonoscopy techniques and their prospective implications for multimodal medical applications. With this goal, we begin by assessing the current data-centric and model-centric landscapes through four tasks for colonoscopic scene perception, including classification, detection, segmentation, and vision-language understanding. Our assessment reveals domain-specific challenges and underscores the need for further multimodal research in colonoscopy. To address these gaps, we establish three foundational initiatives: a large-scale multimodal instruction tuning dataset ColonINST, a colonoscopy-designed multimodal language model ColonGPT, and a multimodal benchmark. To facilitate continuous advancements in this rapidly evolving field, we provide a public website for the latest updates: https://github.com/ai4colonoscopy/IntelliScope.
@@
Colonoscopy, Survey, Vision-language, Multimodal language model, Multimodal benchmark, Abdomen, Healthcare AI.
I Introduction
Despite declining colorectal cancer (CRC) rates in high-income countries, it remains the third most diagnosed cancer worldwide and is increasing in developing countries [1]. As an efficient method for CRC screening, colonoscopy utilises a flexible camera-equipped tube to visually examine the colon’s interior. As illustrated in Fig. (1-a), this clinical procedure also facilitates intervention with specialised instruments such as snares, forceps, and cautery devices to remove precancerous growths, such as serrated and adenomatous polyps. A recent study [2] indicates that incorporating artificial intelligence (AI) into colonoscopy reduces the miss rate of colorectal neoplasia by approximately 50% compared to traditional methods. This success motivates us to investigate the frontiers in intelligent colonoscopy.
Colonoscopy, an endoscopic optical imaging technique, usually presents challenges such as non-uniform illumination and homogeneity of visual patterns, that differ from those of general-purpose imaging data, e.g., ImageNet [3], due to the complex and folded anatomy of the colon. This suggests that special methods are needed to interpret the colonoscopic data. In response, we begin with an investigation of the latest intelligent techniques for colonoscopy, assessing the current landscape to sort out domain-unique challenges and underexplored areas. Our analysis reveals that multimodal research in colonoscopy remains largely untapped. To bridge this gap, we contribute the following three efforts to the community, as illustrated in Fig. (1-b).
Figure 1: Overview of colonoscopy and study highlights: We depict (a) the anatomy of the large intestine (colon) within the digestive tract, the polypectomy procedure during colonoscopy examination, and the colonoscope components. (b) summarises three highlights of this study.
@@ I Introduction
Contribution. (a) We investigate the latest research progress in four colonoscopic scene perception tasks (refer to Fig. 2) from both data-centric and model-centric perspectives. Our investigation summarises key features of 63 datasets and 137 representative deep learning techniques published since 2015. Additionally, we highlight emerging trends and opportunities for future study. (b) We introduce ColonINST, a pioneering instruction tuning dataset tailored for multimodal research, aimed at instructing models to execute user-driven tasks interactively. Assembled from 19 publicly available sources, the ColonINST dataset contains 303,001 colonoscopy images across 62 sub-categories, reflecting diverse scenarios encountered in colonoscopy procedures. We expand these visual samples in two aspects. First, we leverage the multimodal AI chatbot, GPT-4V [4], to generate 128,620 medical captions. Second, we restructure 450,724 human-machine conversations for multimodal adaptation. (c) Leveraging the instruction tuning data, we build a multimodal language model, ColonGPT  to assist endoscopists through interactive dialogues. To ensure reproducibility for average community users, we implement ColonGPT in a resource-friendly way, using a 0.4B-parameter visual encoder SigLIP-SO [5] and a 1.3B-parameter lightweight language model Phi1.5 [6]. Unlike previous vision-language (VL) linking methods [7, 8, 9] that employ multilayer perceptrons to handle all tokens equally from the visual encoder, we propose a multigranularity adapter to selectively sample visual tokens according to their significance. This strategy reduces the visual tokens to $\sim$34of the original number without compromising performance, securing the top spot in our newly created multimodal benchmark across three tasks. Importantly, our ColonGPT can be trained within seven hours on two NVIDIA H200 GPUs, facilitating rapid proof-of-concept development for subsequent research.
@@
Scope. This study differs from previous works in several aspects. Earlier surveys on traditional [10] and deep learning [11, 12, 13, 14] methods conducted before 2020 are now out of date. Although a recent study [15] explores various applications of colonoscopy, such as quality analysis and anomaly detection, it lacks numerical validation. Other benchmarks [16, 17, 18] are limited to specific narrow research subfields. By contrast, we delve into four tasks related to colonoscopic scene perception and evaluate their current state to identify key challenges and under-researched areas. Importantly, our vision goes beyond by laying the foundations for the upcoming advancements in the multimodal domain. To this end, we undertake three initiatives to community: a multimodal instruction tuning dataset, a multimodal language model, and a multimodal benchmark.
@@
Organisation. The remaining sections are structured as follows: §II provides a historical background and discusses the unique challenges of the domain. In §III, we investigate 63 datasets related to colonoscopy, followed by a survey of 137 deep models in §IV. In §V, we introduce three initiatives towards the multimodal era: the creation of ColonINST dataset, the technical details of ColonGPT model, and a comparative multimodal benchmark along with ablative analyses. Finally, this paper is concluded in §VI.
@@ C-E Crohn’s Disease [53] FMOLB’21 467 164 #C1 $\rightarrow$ Crohn’s lesions -
II Background
II-A Origin and evolution
The history of colonoscopy has two key milestones. In 1968, gastrointestinal surgeons Hiromi Shinya and William Wolff found a link between colonic polyps and intestinal tumours, but they lacked equipment to examine them. In 1969, they discovered Corning Incorporated’s optical fibres and collaborated with Olympus to create the fiberoptic colonoscope, a groundbreaking device to examine the colon and remove polyps using wire loops. The second milestone came in 1983 with the introduction of the electronic colonoscope [79], which allows visualisation of the colon on a screen and polyp removal using a polypectomy snare, enhancing detection rates and reducing bleeding. The 21st century brings the AI era, where computer-aided diagnosis systems provide greater precision and reliability in procedures [80]. This study explores the transformative impact of intelligent techniques for colonoscopy, a type of endoscopy [81], while other related areas such as laparoscopy [82] are discussed in appendix.
II-B Intrinsic traits and domain-unique challenges
@@
III Revisiting Colonoscopy Data
III-A Medical data for colonoscopy
@@
$\bullet$ Detection datasets provide both categorical and localisation labels for targets of interest, such as colonic diseases [54, 57, 62], accessory instruments [42, 34, 42], polyps [48, 59, 23, 64, 72], endoscopic artefacts [45, 36], and gastrointestinal diseases [56, 43]. In addition, the organisation of competitions has accelerated growth within the colonoscopy community by establishing shared platforms for data collection and model evaluation, greatly advancing research areas like the detection of intestinal disease [40] and polyp [61, 69, 38].
@@ III-B Discussion
$\bullet$ Data granularity requires improvement to better understand patient conditions and treatment efficacy. (a) More than a quarter of polyp-containing datasets provide fine-grained categorisation, often with inadequate detail. For example, BKAI-Small/-Large [46] provide two instance annotations of neoplastic and non-neoplastic polyps. ColonoscopicDS [26] categorises at the video level into hyperplasic, serrated, and adenoma lesions. SUN-database [54] provides fine-grained labels, documenting measurements of polyp size (height, width) and morphology (pedunculated, sessile, flat), along with their anatomical locations (e.g., rectum, sigmoid colon). Several data-centric areas remain under-explored, such as temporal lesion evolution recording, granularity improvement, graded severity tagging, and instance-level target annotation. (b) Label orthogonality, an often overlooked issue, treats classes as isolated entities. Current works seldom discuss potential inter-class correlations, such as the co-occurrence of inflammatory bowel disease with erosion symptoms, Crohn’s disease with fistula complications, or colorectal cancer accompanied by bleeding. Future studies should consider causality [88] and comorbidity [89] to address these correlations effectively.
@@
TABLE II: Summary of classification models in colonoscopy. Dataset: CU=CU-ColonDB [25], CDS=ColonoscopicDS [26], Private= private data, HK=HyperKvasir [43], KC=Kvasir-Capsule [56]. Backbone: CaffeNet [97], D-121=DenseNet121 [98], R-12/-18/-50/ -101=ResNet12/18/50/101 [99], ViT-S16/ViT-B16 [100], MobV2=MobileNetV2 [101], R50-Att=ResNet50 with attention module [102], C3D [103], Inc-v3=Inceptionv3 [104], I3D [105]. “Customised” means a base network modified for the current task or a model independent of the base network choice. Head: classifier implemented by the fully connected (FC) and support vector machine (SVM) layers, or using the $\ell^{2}$ norm to measure the disparity between the input and output. Arch: the architectures shown in Fig. 3. Sup: learning strategies such as fully supervised (FS), semi-supervised (SS), unsupervised (US), and weakly supervised (WS). For simplicity, the following tables use consistent abbreviations unless specified otherwise.
@@
IV Revisiting Colonoscopy Models
In this section, we investigate 137 deep learning models for colonoscopic scene perception, sourced from leading conferences or journals published since 2015. First, three widely recognised topics are described, including 18 classification models (§IV-A), 24 detection models (§IV-B), and 86 segmentation models (§IV-C). Their architectural designs are classified into five subtypes as presented in Fig. 3. Lastly, we discuss nine VL-related models in §IV-D.
IV-A Classification models
@@
V2I-DETR [148] BIBM’24 video-to-image knowledge distillation SUN SUN R-50 OS BF#2 FS -
@@
$\bullet$ Remarks. We observe three aspects of the above classification models. (a) Novel visual backbones, like the state space model [149], remain underexplored. Furthermore, reformulating the classification paradigm within VL models, e.g., CLIP [150], can yield unexpected results. (b) Benchmarks for multicategory classification remain underexplored. The Kvasir series [28, 43, 56] offers valuable sources for further study. We will explore these public data on their potential synergy in §V-A. (c) Several new task settings have emerged in colonoscopy. For example, Tian et al.[59] recognise abnormal frames from colonoscopy videos from an out-of-distribution view. DAFON [115] solve an open-set classification problem within a few-shot framework.
IV-B Detection models
@@
U-KAN [239] AAAI’25 U-shaped Kolmogorov-Arnold network [240] C6 C6 Customised MF#1 - FS Link
@@
TABLE V: Comparison of image polyp segmentation models. Models are evaluated using the mean scores (%) of structure measure ($\mathcal{S}$ [258]) and Dice coefficient ($\mathcal{D}$) on two test sets, with boxplots illustrating the distribution of their consistency and variability across test cases. The rankings for each model are marked in grey colour.
@@
IV-C Segmentation models
Compared to the above two topics, the segmentation research appears to be well-established (see Tab. IV).
@@ IV-C Segmentation models
[…]
Figure 4: Details of ColonINST. (a) Three sequential steps to create the instruction tuning dataset for multimodal research. (b) Numbers of colonoscopy images designated for training, validation, and testing purposes. (c) Data taxonomy of three-level categories. (d) A word cloud of the category distribution by name size. (e) Caption generation pipeline using the VL prompting mode of GPT-4V [4]. (f) Numbers of human-machine dialogues created for four tasks.
@@
$\bullet$ Remarks. To reflect current field progress, we analysed 14 open-source image segmentation models on two popular test datasets, as shown in Tab. V. All models are trained on the same dataset from Fan et al.’s benchmark [175]. First, we observe that current learning strategies are underexplored, as evidenced by BoxPolyp, a weakly supervised model that obtains the best $\mathcal{D}$ score (91.84%) on Kvasir-SEG. In addition, some models achieve better performance, yet exhibit wider interquartile ranges in boxplots, indicating their variability in predictions. For example, Polyp-PVT, which ranks highest in the $\mathcal{S}$ score on Kvasir-SEG, shows a wider interquartile range than other models like SSFormer. From these results, we suggest several promising opportunities for future study. (a) The current gold benchmark [175] comprises less than 1.5K samples and is focused on a category (polyp). In general, scaling up both the data size and diversity could be a natural way to improve robustness and generalisability. This demand is driving innovations, such as developing a semi-auto image annotator [166, 165] to reduce expert labour and synthesising high-fidelity content via diffusion [261] and autoregressive [262] techniques. (b) Moreover, infinite data scaling is not sustainable. Developing data-efficient strategies [94, 95, 96] that require fewer or weaker labels is more cost-effective for average users in the community. (c) Finally, providing procedural support to physicians is essential, including anomaly detection, navigation planning, risk assessment, and intervention advice. We can adopt innovations from similar topics [263].
IV-D VL models